OpenStudy (anonymous):

Abigail orders a large 2 topping pizza for $14 and a medium 1 topping pizza for $10. If all costs of a medium pizza are 80% of a similar large pizza, find the price of a large plain cheese pizza and the cost of each large pizza topping.

OpenStudy (retireed):

Okay, kind of confusing to me. Lots of data and I am having trouble with an equation. Let's see what we know.... Large 2 top = 14 Medium 1 top = 10 Medium X top = 0.80 Large X top What is the price of a Large No top? What is price of a large 1 top? Correct?

OpenStudy (anonymous):

Yes! I am having a problem coming up with an equation.

OpenStudy (retireed):

We know a Medium 2 Top will equal 80% of a Large 2 Top... Medium 2 Top = 0.08 (14) = 11.2 Correct ? From here if we subtract Medium 2 Top from Medium 1 top we can determine the price of and medium single topping.

OpenStudy (anonymous):

Looks right. So what do I subtract from 11.20?

OpenStudy (retireed):

Medium 1 topping. It is given in the question.

OpenStudy (anonymous):

Ok so thats 1.20 per topping! So that helps us find the cost of a medium topping. So to find the cost of a large I would have to subtract 0.80?

OpenStudy (retireed):

Correct, $1.20 is the cost of a medium topping. I am confused as to how you came up subtracting $0.80 to find the cost of a large, since the cost of a large topping would be greater than $1.20. Can you clarify your reasoning?

OpenStudy (anonymous):

Sorry I mean adding. So the price of a large topping would be 2.00 and the cost of a large cheese pizza would be $10?

OpenStudy (retireed):

I'm checking your math. How did you get $2 for the cost of a large topping? I'm getting something else, but I could be wrong.

OpenStudy (anonymous):

If a medium topping is 1.20, a medium topping is .80 of a large topping so a large topping would be $2?

OpenStudy (retireed):

Okay, great that is the same reasoning I used, but I think our math is incorrect. This is what I did and see if you agree.... MT1 is 1 medium topping LT1 is one large topping MT1 = 0.8 LT1 the original ratio of medium to large and we know MT1 is $1.20 So LT1 = MT1 / 0.8 = 1.2 / 0.8 ???? What do you get ?

OpenStudy (retireed):

Here is a quasi equation form of my reasoning L2 =14 cost of a large 2 topping pizza M1=10 cost of a medium 1 topping pizza M = 0.8L ratio between medium and large 0.8L2 -M1 = MT1 cost of one medium topping MT1 = 0.8 LT1 going back to the original ratio MT1 / 0.8 = LT1 cost of one large topping LT1 is a critical value if you know the cost of a large topping you should be able to find the price or cost of a L1 and a large no topping cheese pizza